solve for x .
log(x+7)-logx=3

Answer by stanbon(48535)

log(x+7)-logx=3
----------------
log [(x+7)/x] = 3
(x+7)/x = 10^3
x+5 = 1000x
999x = 5
x = 0.0050505....
===========
Cheers,
Stan H.